OpenStudy (anonymous):

PLEASE HELP ME Solve, using linear combination. 5x + 2y = -4 -3x + 2y = 12 A. (-2,3) B. (0,-2) C. (2, -7) D. (4,0)

OpenStudy (anonymous):

@triciaal @iambatman @One098 @pooja195 @Abhisar @ShadowLegendX @Directrix @Compassionate @bibby @mathmath333

OpenStudy (one098):

Plug in each of them until one of the pairs solves the linear equations.

OpenStudy (triciaal):

@One98 and if not given the choices?

OpenStudy (anonymous):

Try each one until you find it. it is really easy

OpenStudy (anonymous):

A

OpenStudy (one098):

@triciaal Yeah, I know what you mean, but I was just telling her that since she does have choices, you just plug them in.

OpenStudy (anonymous):

I got A

OpenStudy (one098):

@JimStanSGC Leave.

OpenStudy (triciaal):

@laylaflower to solve by combination the object is to make the coefficients of one variable be equal and opposite so they will cancel out when added. this will give one equation with one unknown variable. solve then substitute in either of the originals to solve for the other variable.

OpenStudy (anonymous):

Your correct A

OpenStudy (anonymous):

thank you all ! :)

OpenStudy (anonymous):

Your welcome

OpenStudy (aaronandyson):

you can multiply the first equation by 2 and get rid of variable y by adding both equations together you get 12x-4y = 10 5x + 4y = 24 17x = 34 x must be 2 if x = 2 then you plug into the equation 12 - 2y = 5 7 = 2y y must be 3.5 i think this is right ..........

OpenStudy (triciaal):

for this one just subtract since the y's are equal solve for x 8 x = -16 so x = -2

OpenStudy (anonymous):

http://openstudy.com/study#/updates/54b2a7dbe4b0dbc68afeefd1 new question